OpenStudy (anonymous):

PLEASE CAN ANYONE HELP.....Mark's Store sells cashews for $5.00 per pound and peanuts for $1.50 per pound. Mark decides to mix 30 pounds of peanuts with some cashews and sell the mixture for $3 per pound. How many pounds of cashews should be mixed with the peanuts so that the mixture will produce the same revenue as would sell the nuts separately?

OpenStudy (anonymous):

Solve the following equations for c and p, the pound weight of cashews and peanuts respectively. Dollars are expressed as pennies. {c + p = 30, 500 c + 150 p = 300*30}